OpenStudy (anonymous):

given f(x)=-3x+4 and g(x)=x+7, what is the value of g(f(2))..explain it to me please..

OpenStudy (anonymous):

f(2) = -3(2) + 4 =2 Now put 2 into g(x) = 2 + 7 = 9

OpenStudy (anonymous):

gf(x) means 'the value of f at x substituted into the function g'.

OpenStudy (anonymous):

sorry - f(2) = -2 and gf(2) = -2+7 = 5

OpenStudy (anonymous):

g(f(2))=(-3x+4)+7 =-3x+4+7 = -3x+11 subtitution x=2 = -3(2)+11 = -6+11 = 5

OpenStudy (anonymous):

If I deserve medal then give me one.

OpenStudy (anonymous):

\[f(2) = -3\times 2+4=-6+4=-2\] \[g(-2)=-2+7=5\] Nancy Lam got it!